# Comparative Microscopic Study of Entropies and their Production

### v1

@@ 1 Motivation
In view of the pivotal roles of entropy, temperature and the second law in science, it is surprising that more than 150 years after their inception [1] scientists have not even approximately agreed on their basic microscopic definitions. Instead, different schools of thought have formed that differ widely both from a philosophical-conceptual and (what will be our focus here) quantitative-numerical point of view. In addition, the topic appears unnecessarily mystified as evidenced, for instance, by von Neumann’s often quoted statement “no one knows what entropy really is” (made while discussing with Shannon about a name for his “entropy” concept [2]). This quote seems unfortunate as historical evidence indicates that von Neumann had a clear opinion on which microscopic entropy definition to use in statistical physics [3, 4, 5], which—to add to the confusion—is not what is nowadays known as von Neumann entropy.
@@
With advances in computer technologies it has now become possible to simulate isolated quantum (and also classical) systems, which are large enough to exhibit thermodynamic behaviour, without approximations. Consequently, a number of interesting case studies emerged that focused on the nonequilibrium dynamics of thermodynamic entropy from first principles [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. [^2] The literature is even wider if one includes equilibrium systems or effective dynamics such as master equations. This is not our focus here because both cases potentially mask a large part of the problem. However, somewhat echoing the criticism above, a study critically comparing a variety of entropy definitions on the same footing is not known to us.
@@ 2 Framework
Our idea is to study a paradigmatic nonequilibrium process, namely the flow of heat between two bodies $A$ and $B$ (similar but simpler versions of the model below have been studied, e.g., in Refs. [45, 46]). Each system $X\in\{A,B\}$ is modeled with a Hilbert space ${\cal{H}}_{X}$ with dimension $D_{X}$ and Hamiltonian $H_{X}=\sum_{j=1}^{D_{X}}{\epsilon}_{j}^{X}|{\epsilon}_{j}{\rangle\langle}{\epsilon}_{j}|_{X}$.
@@
To model two large bodies that interact sufficiently strongly to allow efficient energy exchanges while at the same time being weakly coupled such that the sum of the local energies is approximately conserved, we need to find a delicate balance between $\delta_{V}$ (which determines the sparsity of the Hamiltonian) and the coupling strength $\lambda$. What we found to work well is the choice
$\lambda^{2}=\frac{1}{D^{A}_{\text{eff}}D^{B}_{\text{eff}}}\left(\frac{1}{5}\frac{\Delta_{E}^{2}}{\delta V}\right)^{2},~{}~{}~{}\delta_{V}=\frac{\delta_{E}}{6},$ (1)
where $\Delta_{E}=E_{b}-E_{a}$ (not to be mixed up with the coarse graining width $\delta_{E}$). Moreover, in all numerical results below we set $E_{\text{min}}=0$, $E_{\text{max}}=2.1$, $E_{a}=0.42$, $E_{b}=1.4$ and $\delta_{E}=0.14$. This implies $M=15$ coarse grained energy windows in the full interval $[E_{\text{min}},E_{\text{max}}]$ and seven windows in the dynamically relevant subinterval $[E_{a},E_{b}]$.
@@ 3.1 Class A: Two normal systems
Turning to the comparison of all eleven different entropies in Fig. 3(e), we see that all of them tend to increase, i.e., they satisfy the second law of thermodynamics. However, both the Boltzmann surface and Gibbs volume entropies $S_{B}({\langle{E}\rangle})$, ${\langle{S_{B}}\rangle}$, $S_{G}({\langle{E}\rangle})$ and ${\langle{S_{G}}\rangle}$ stay almost constant. $S_{B}({\langle{E}\rangle})$ and $S_{G}({\langle{E}\rangle})$ show even a short upward and downward fluctuation related to the discrete coarse graining. Also the canonical entropy production is significantly smaller than the production of the entanglement, diagonal and cgo entropies. It is further noticeable that the diagonal entropy $S_{\text{diag}}$ (for IC2 and IC3) almost perfectly matches the local cgo entropy $S_{\text{cgo}}^{\otimes}$, while the entanglement entropy $S_{\text{ent}}$ (for IC3) evolves in parallel but not very close to the cgo entropy $S_{\text{cgo}}$. Moreover, as it should be for normal systems, all volume-related entropies show the same behaviour, just shifted upwards to larger numerical values (that is, the entropy productions are the same if we interchange the letters $B$ by $G$).
Finally, it is worth to notice that the cgo entropy $S_{\text{cgo}}$ does not satisfy von Neumann’s $H$-theorem [3, 4] (see also Refs. [85, 86, 87]), i.e., it does not saturate to its predicted maximum value $\ln(D_{\text{eff}}^{A}D_{\text{eff}}^{B})$, which is indicated by the thin horizontal gray line. We attribute this discrepancy to the fact that the state is not confined to a narrow microcanonical energy shell, i.e., we believe the DOS of the global Hamiltonian varies significantly over the dynamically relevant energy range (we can not prove this fact as an exact diagonalization of the total Hamiltonian is out of reach). In this case, non-negligible correlations between the global and local energy eigenbasis could jeopardize the applicability of von Neumann’s $H$-theorem.
@@ 3.3 Class C: Normal system coupled to a negative heat capacity system
Finally, we consider a normal system $A$ with DOS $\mu_{A}(\epsilon)=\exp(\sqrt{10\epsilon})$ coupled to a system $B$ with DOS $\mu_{B}(\epsilon)=\Theta(E_{b}-\epsilon)\exp[(3\epsilon/2)^{2}]$. Since $C\sim-\partial_{\epsilon}^{2}\ln\mu_{B}(\epsilon)=-3/2$ this describes a system with negative heat capacity. Moreover, the Heaviside step function $\Theta(x)$ introduces a cutoff outside the dynamically relevant regime (physically, this could be justified by a gaped DOS or an additional conservation law). The resulting situation is sketched in Fig. 2(C). The initial macrostate is $p_{\alpha,\beta}(0)=\delta_{\alpha,1}\delta_{\beta,7}$ and the effective dimensions are $D_{\text{eff}}^{A}=372$ and $D_{\text{eff}}^{B}=392$. All numerical results are displayed in Fig. 5 with the same conventions as in Figs. 3 and 4.
@@ 4 Summary
    Cgo entropies, whether of surface or volume type, show a strong increase similar to the entanglement or diagonal entropy. This highlights that probabilistic considerations, captured by a Shannon entropy term in cgo entropies, seem important for the validity of the second law. In particular, we repeat that these probabilities are genuinely of quantum origin in our model: they do not arise from classical ensemble averages or a subjective lack of classical knowledge. Finally, despite its increase, the cgo entropy $S_{\text{cgo}}$ does not satisfy von Neumann’s $H$-theorem, presumably because of the too widespread energy interval considered here.

### v2

@@ 1 Motivation
In view of the pivotal roles of entropy, temperature and the second law in science, it is surprising that more than 150 years after their inception [1] scientists have not even approximately agreed on their basic microscopic definitions. Instead, different schools of thought have formed that differ widely both from a philosophical-conceptual and (what will be our focus here) quantitative-numerical point of view. In addition, the topic appears unnecessarily mystified as evidenced, for instance, by the often quoted statement “no one knows what entropy really is”allegedly made by von Neumann while discussing with Shannon about a name for his “entropy” concept [2], even though it is unclear whether von Neumann truly said it [3]. This quote also seems particularly unfortunate as historical evidence indicates that von Neumann had a clear opinion on which microscopic entropy definition to use in statistical physics [4, 5, 6], which—to add to the confusion—is not what is nowadays known as von Neumann entropy.
@@
With advances in computer technologies it has now become possible to simulate isolated quantum (and also classical) systems, which are large enough to exhibit thermodynamic behaviour, without approximations. Consequently, a number of interesting case studies emerged that focused on the nonequilibrium dynamics of thermodynamic entropy from first principles [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. [^2] The literature is even wider if one includes equilibrium systems or effective dynamics such as master equations. This is not our focus here because both cases potentially mask a large part of the problem. However, somewhat echoing the criticism above, a study critically comparing a variety of entropy definitions on the same footing is not known to us.
@@ 2 Framework
Our idea is to study a paradigmatic nonequilibrium process, namely the flow of heat between two bodies $A$ and $B$ (similar but simpler versions of the model below have been studied, e.g., in Refs. [51, 52, 53]). Each system $X\in\{A,B\}$ is modeled with a Hilbert space ${\cal{H}}_{X}$ with dimension $D_{X}$ and Hamiltonian $H_{X}=\sum_{j=1}^{D_{X}}{\epsilon}_{j}^{X}|{\epsilon}_{j}{\rangle\langle}{\epsilon}_{j}|_{X}$.
@@
In all numerical results below we set $E_{\text{min}}=0$, $E_{\text{max}}=2.1$, $E_{a}=0.42$, $E_{b}=1.4$ and $\delta_{E}=0.14$. This implies $M=15$ coarse grained energy windows in the full interval $[E_{\text{min}},E_{\text{max}}]$ and seven windows in the dynamically relevant subinterval $[E_{a},E_{b}]$. Moreover, the coupling strength $\lambda$ and the bandedness $\delta_{V}$ are chosen to satisfy the following demands:
  1. 1.
    We want a weak coupling between $A$ and $B$ such that ${\langle{H_{A}+H_{B}}\rangle}(t)\approx{\langle{H_{A}+H_{B}}\rangle}(0)$. In this regime bulk properties dominate surface properties and it becomes meaningful to talk about thermodynamic properties (energy, entropy, temperature) of $A$ or $B$ alone. Specifically, we define weak coupling by demanding that the fluctuations in the interaction energy ${\langle{V^{2}}\rangle}_{\text{mic}}$, where ${\langle{\dots}\rangle}_{\text{mic}}$ denotes a microcanonical average, are small compared to ${\langle{H_{A}+H_{B}}\rangle}^{2}_{\text{mic}}$. Using the bandedness of $V$, we estimate ${\langle{V^{2}}\rangle}_{\text{mic}}\approx\lambda^{2}D_{\text{eff}}^{2}\delta_{V}/\Delta_{E}$, where $\Delta_{E}=E_{b}-E_{a}$ denotes the energy range of the dynamically relevant regime (not to be mixed up with the coarse graining width $\delta_{E}$). Moreover, we have ${\langle{H_{A}+H_{B}}\rangle}^{2}_{\text{mic}}\approx 4\Delta_{E}^{2}$, tacitly assuming the arbitrary energy offset $E_{a}$ to be subtracted.
  2. 2.
    Opposite to the weak coupling requirement, we also need to demand that the energy levels of $A$ and $B$ are sufficiently strongly coupled to allow efficient energy exchanges. This implies that the coupling strength $\lambda$, which broadens the levels in an interacting system, should exceed the mean level spacing $\Delta_{E}/D_{\text{eff}}$ of $A$ or $B$ (this is a rough approximation assuming that all $D_{\text{eff}}$ levels are distributed evenly in $[E_{a},E_{b}]$).
  3. 3.
    Finally, we need to find a compromise between a very sparse matrix with small $\delta_{V}$, which allows to simulate systems with larger $D_{\text{eff}}$, and a large $\delta_{V}$ such that the eigenstates of the global Hamiltonian $H_{A}+H_{B}+V$ are sufficiently delocalized with respect to the local energy eigenbasis of $H_{A}+H_{B}$: if $\delta_{V}$ is too small, energy transport becomes blocked due to Anderson localization.
What we found to work well is the choice
$\lambda^{2}=\frac{1}{D^{A}_{\text{eff}}D^{B}_{\text{eff}}}\left(\frac{1}{5}\frac{\Delta_{E}^{2}}{\delta_{V}}\right)^{2},~{}~{}~{}\delta_{V}=\frac{\delta_{E}}{6}.$ (1)
@@
Finally, all entropy notions are summarized in Table 1. Note that they are all based on the common idea to treat the system in a coarse way using incomplete information, just how they do it precisely differs from definition to definition. In contrast, the full microscopic von Neumann entropy $H_{\text{vN}}[\psi(t)]$ is zero and remains constant all the time.
name symbol definition comment
Boltzmann entropy $S_{B}({\langle{E}\rangle})$ $\ln W_{{\langle{\alpha}\rangle}}+\ln W_{{\langle{\beta}\rangle}}$ related to $\beta_{B}$ in the
of the average continuum limit $\delta_{E}\rightarrow 0$
averaged ${\langle{S_{B}}\rangle}$ ${\langle{\ln W_{\alpha}}\rangle}+{\langle{\ln W_{\beta}}\rangle}$
Boltzmann entropy
cgo entropy $S_{\text{cgo}}$ $H_{\text{Sh}}[p_{\alpha,\beta}]+{\langle{S_{B}}\rangle}$
local cgo entropy $S_{\text{cgo}}^{\otimes}$ $H_{\text{Sh}}[p_{\alpha}\otimes p_{\beta}]+{\langle{S_{B}}\rangle}$
volume entropy $S_{G}({\langle{E}\rangle})$ $\ln\Omega_{{\langle{\alpha}\rangle}}+\ln\Omega_{{\langle{\beta}\rangle}}$ related to $\beta_{G}$ in the
of the average continuum limit $\delta_{E}\rightarrow 0$
averaged ${\langle{S_{G}}\rangle}$ ${\langle{\ln\Omega_{\alpha}}\rangle}+{\langle{\ln\Omega_{\beta}}\rangle}$
volume entropy
cgo entropy $S_{G,\text{cgo}}$ $H_{\text{Sh}}[p_{\alpha,\beta}]+{\langle{S_{G}}\rangle}$
(volume version)
local cgo entropy $S_{G,\text{cgo}}^{\otimes}$ $H_{\text{Sh}}[p_{\alpha}\otimes p_{\beta}]+{\langle{S_{G}}\rangle}$
(volume version)
canonical entropy $S_{\text{can}}({\langle{E}\rangle})$ $H_{\text{vN}}[\pi_{A}]+H_{\text{vN}}[\pi_{B}]$ related to $\beta_{\text{can}}$
entanglement $S_{\text{ent}}$ $H_{\text{vN}}[\rho_{A}]+H_{\text{vN}}[\rho_{B}]$
entropy
diagonal entropy $S_{\text{diag}}$ $H_{\text{Sh}}[p_{j}^{A}]+H_{\text{Sh}}[p_{j}^{B}]$
Table 1: Overview of all the different entropy notions compared in this work. We dropped the dependence on time $t$ for notational simplicity.
@@ 3 Numerical results
Moreover, all numerical results are displayed for a single choice of the random matrix Hamiltonian and a single choice for the three types of initial states. We do not perform any averages over them, but we have observed that the coarse behaviour is quite insensitive to the precise choice of Hamiltonian or initial state (not displayed here).
@@ 3.1 Class A: Two normal systems
Turning to the comparison of all eleven different entropies in Fig. 3(e), we see that all of them tend to increase, i.e., they satisfy the second law of thermodynamics. However, both the Boltzmann surface and Gibbs volume entropies $S_{B}({\langle{E}\rangle})$, ${\langle{S_{B}}\rangle}$, $S_{G}({\langle{E}\rangle})$ and ${\langle{S_{G}}\rangle}$ stay almost constant. The upward and downward jumps of $S_{B}({\langle{E}\rangle})$ and $S_{G}({\langle{E}\rangle})$ are caused by a brief transition of the average energy from one discrete energy window to another. Also the canonical entropy production is significantly smaller than the production of the entanglement, diagonal and cgo entropies. It is further noticeable that the diagonal entropy $S_{\text{diag}}$ (for IC2 and IC3) almost perfectly matches the local cgo entropy $S_{\text{cgo}}^{\otimes}$, while the entanglement entropy $S_{\text{ent}}$ (for IC3) evolves in parallel but not very close to the cgo entropy $S_{\text{cgo}}$. Moreover, as it should be for normal systems, all volume-related entropies show the same behaviour, just shifted upwards to larger numerical values (that is, the entropy productions are the same if we interchange the letters $B$ by $G$).
Finally, it is worth to notice that von Neumann’s $H$-theorem [4, 5] (see also Refs. [92, 93, 94, 95]) establishes the conditions under which the cgo entropy $S_{\text{cgo}}$ saturates to its maximum value $\ln(D_{\text{eff}}^{A}D_{\text{eff}}^{B})$ (thin horizontal gray line), which does not happen in our example. We attribute this discrepancy to the fact that the perturbation $V=\lambda R$ is too banded or, equivalently, the global state is not confined to a sufficiently narrow energy shell. In this case, the eigenfunctions of $H_{A}+H_{B}+V$ are not delocalized over the entire dynamically relevant Hilbert space (we can not prove this fact as an exact diagonalization of the total Hamiltonian is out of reach) such that the resulting non-negligible correlations between the global and local energy eigenbasis could jeopardize the applicability of von Neumann’s $H$-theorem.
@@ 3.3 Class C: Normal system coupled to a negative heat capacity system
Finally, we consider a normal system $A$ with DOS $\mu_{A}(\epsilon)=\exp(\sqrt{10\epsilon})$ coupled to a system $B$ with DOS $\mu_{B}(\epsilon)=\Theta(E_{b}-\epsilon)\exp[(3\epsilon/2)^{2}]$. Since $C\sim-\partial_{\epsilon}^{2}\ln\mu_{B}(\epsilon)=-3/2$ this describes a system with negative heat capacity as it could arise in systems with long range interactions [10, 11, 12, 13, 14, 15]. Moreover, the Heaviside step function $\Theta(x)$ introduces a cutoff outside the dynamically relevant regime (physically, this could be justified by a gaped DOS or an additional conservation law). The resulting situation is sketched in Fig. 2(C). The initial macrostate is $p_{\alpha,\beta}(0)=\delta_{\alpha,1}\delta_{\beta,7}$ and the effective dimensions are $D_{\text{eff}}^{A}=372$ and $D_{\text{eff}}^{B}=392$. All numerical results are displayed in Fig. 5 with the same conventions as in Figs. 3 and 4.
@@ 4 Summary
    Cgo entropies, whether of surface or volume type, show a strong increase similar to the entanglement or diagonal entropy. This highlights that probabilistic considerations, captured by a Shannon entropy term in cgo entropies, seem important for the validity of the second law. In particular, we repeat that these probabilities are genuinely of quantum origin in our model: they do not arise from classical ensemble averages or a subjective lack of classical knowledge. Finally, despite its increase, von Neumann’s $H$-theorem for cgo entropy $S_{\text{cgo}}$ does not apply, presumably because of the too narrow band structure of $V$ (or the too widespread energy interval) considered here.